### v1

# INSIGHT: Bridging the Student-Teacher Gap in Times of Large Language Models

Abstract
The rise of AI, especially Large Language Models, presents challenges and opportunities to integrate such technology into the classroom. AI has the potential to revolutionize education by helping teaching staff with various tasks, such as personalizing their teaching methods, but it also raises concerns, for example, about the degradation of student-teacher interactions and user privacy. This paper introduces INSIGHT, a proof of concept to combine various AI tools to assist teaching staff and students in the process of solving exercises. INSIGHT has a modular design that allows it to be integrated into various higher education courses. We analyze students’ questions to an LLM by extracting keywords, which we use to dynamically build an FAQ from students’ questions and provide new insights for the teaching staff to use for more personalized face-to-face support. Future work could build upon INSIGHT by using the collected data to provide adaptive learning and adjust content based on student progress and learning styles to offer a more interactive and inclusive learning experience.
Keywords:
AI Teaching Assistant Learning Support Student-Teacher Interaction.
1 Introduction
Due to the rising popularity of Artificial Intelligence (AI), various educational tools incorporating AI have been developed in recent years. AI can provide personalized learning experiences by tailoring educational experiences to individual learners’ needs and abilities [4]. For example, adapting content and pace can improve learning effectiveness and efficiency [12, 18]. AI enables personalized learning at any time, making education more efficient and accessible to larger groups [1]. Although most research focuses on K-12 settings, personalized learning is also promising for higher education, which is the main focus of this work, as AI can consider their unique motivations and circumstances [18]. Recent research also focuses on making Large Language Models (LLMs) fit for the classroom by reducing hallucinations, which is achieved by using course materials as input for the answer [13]. In mathematics education, LLMs show promise in developing students’ problem-solving skills and critical thinking abilities by enabling them to engage more deeply with open-ended problems through iterative questioning, clarification of concepts, and exploration of multiple solution paths [14]. Nitze [8] developed a prototype to simulate oral exams in STEM education to reduce educator workload and give students individual feedback in the early stages of their academic journey.
Despite the numerous benefits AI offers to students and teaching staff, its integration into education raises several concerns. First, AI can degrade human interaction in education [2, 10], as it reduces face-to-face engagement and creates excessive dependence on the technology [1, 7], which could lead to worse educational experiences for students. Second, as students start asking more questions to LLMs, the teaching staff might miss important cues, such as gaps in comprehension of the learning material [7, 10]. Third, privacy is an important concern, as AI tools collect a vast amount of data. OpenAI, for example, collects the user’s input content (including files), device information, and location information [9]. Educational applications, when deployed into a classroom, collect additional data that needs to be stored in a way that is both ethical [6] and compliant with applicable laws, such as the EU’s GDPR. Anthology’s BlackBoard, for example, collects, among other data, a student’s solutions, grades, progress on a per-file basis, average hours in the course, and days of inactivity [16].
To address these challenges, we introduce INSIGHT: INtelligent Student and Instructor Guidance through Human-Centered Technology, which implements different AI tools to support student-teacher interaction in a privacy-aware manner. Our contribution is twofold. First, INSIGHT equips teaching staff with data-driven insights on students’ questions and challenges regarding a specific course topic. Based on these insights, they can support each student in a tailored manner and adapt their course material based on the needs of students. Second, INSIGHT provides students access to an LLM in a monitored environment while ensuring data privacy through explicit opt-in data collection.
2 INSIGHT’s modular framework for AI-powered student-teacher interaction
INSIGHT’s components are designed modularly, enabling INSIGHT to be deployed across various domains without changing its core implementation. We achieve this flexibility by dividing the solution into multiple independently adjustable components, as illustrated in Fig. 1. We have a central system, INSIGHT Core, with a keyword extraction and sentence similarity model and a database. This database contains all the course information and data on student interactions with INSIGHT while solving exercises (e.g., questions to the LLM, FAQ usage, and difficulty ratings). INSIGHT Core is connected to a local LLM and a user interface that can be independently changed based on specific needs. As a proof of concept, we incorporate data from Hasselt University’s Algorithms and data structures course [15] to demonstrate and assess INSIGHT’s feasibility.
Figure 1: INSIGHT has a central system, INSIGHT Core, with a keyword extraction and sentence similarity model and a database containing all course information and student interactions with INSIGHT while solving exercises. INSIGHT Core is connected to a local LLM and a user interface that can be independently changed based on specific needs.
Answering student questions through LLMs.
LLMs have become increasingly widespread, which presents a valuable opportunity for the educational sector to provide frameworks that guide students in making the most of these technologies in a meaningful, ethical, and educationally responsible manner. We use an LLM as the main interaction method between students and INSIGHT. We later analyze students’ interactions with the LLM to provide the teaching staff with data they can use to offer better face-to-face support and increase student-teacher interaction. In this prototype, we used Llama 3.2 3B as it has an above-average output speed and below-average latency, ensuring quick responses [3]. This comes at the cost of intelligence, but in our initial testing, it seems adequate for the use case. The small amount of parameters also makes it easier to deploy on consumer hardware.
Dynamically generated FAQ from student questions.
Student questions to the LLM can help teaching staff identify challenging topics and knowledge gaps. Recurrent questions can be added to an FAQ, supporting students facing the same challenges while informing the teaching staff of a possible knowledge gap. This new information also allows teaching staff to provide better face-to-face support for specific students or groups. To group similar questions in the FAQ, INSIGHT uses the vector embeddings from the all-MiniLM-L6-v2 model to determine similarity. This model was chosen because it was trained on multiple datasets, making it very generalizable, while it maintains a relatively high accuracy and speed [11]. We note that this approach was developed for the given algorithms and data structures and might require fine-tuning to apply it to other courses. Another benefit of grouping similar questions is the ability to return the same answer to these questions. INSIGHT caches the LLM’s answer and provides the teaching staff the ability to edit the answer to their liking, ensuring human-in-the-loop involvement and reducing the variability of LLM answers.
Leveraging student interaction data for teaching insights.
When students engage with an LLM, teaching staff may lose sight of their understanding of the course material, making it harder to provide support and refine the learning resources. To address this, we propose a system that collects and analyzes students’ questions to the LLM, their FAQ usage, and explicit input in the form of difficulty ratings for the exercises they complete. This data provides valuable insights that, over time, can help refine course materials to better align with student needs and expectations. To determine the topics from students’ questions to the LLM, INSIGHT must first identify the topics covered in the exercises. To achieve this, we use keyword extraction to infer the exercise topics from the exercise descriptions. The keyword extraction is a mixed-initiative process that first uses KeyBERT [5] to extract an initial list of keywords from the exercise text. After extraction, the teaching staff can manually review the keywords for each exercise to best represent the tasks and make them suitable for identifying question topics. An example is illustrated in Fig. 2.
Figure 2: The keyword selection is a mixed-initiative process where the initial list of keywords is determined through a keyword extraction algorithm. The teaching staff can further refine the list for final use. An example of an original exercise description for Exercise 1 is as follows: “According to the documentation, which kind of functions do the C++ Standard Library and Java provide to search a value in an array? What is available in the Boost C++ libraries?”
Data anonymization policy.
While it can be helpful for teaching staff to identify exercises and topics students find challenging, not all students might be comfortable sharing identifying data. To address this, students can select anonymous mode, allowing teaching staff access to the data without the students’ names attached. Additionally, INSIGHT uses Ollama to run an LLM locally, ensuring student queries, potentially containing sensitive information, remain private and are never shared with third parties. This method ensures that personal data never leaves the students’ devices without their explicit consent.
User interface.
The teaching staff’s user interface, illustrated in Fig. 3, has two main components: a dynamic FAQ and two visualizations. The dynamic FAQ displays frequently asked student questions, and staff can edit the LLM-generated answers to ensure accuracy. The two graphs visualize FAQ view frequencies and the frequency of course topics in student queries to the LLM. The student interface, illustrated in Fig. 4, also includes the FAQ and an additional chat interface for the LLM.
Figure 3: User interface for the teaching staff, with the dynamic FAQ and two visualizations. The teaching staff can edit the FAQ items and their content. The visualizations show the frequency with which each FAQ item has been viewed and the frequency of topics in the questions from students to the LLM.
Figure 4: The student interface includes the FAQ and a chat window for the LLM. FAQ views are counted, optionally attaching the student’s name with their explicit consent. Chat questions are analyzed to visualize question topics for teaching staff. Students can also rate exercise difficulty to provide explicit feedback.
3 Discussion
INSIGHT presents an approach to integrating AI in education to enhance student-teacher interaction while maintaining transparency in data collection and ensuring users retain control over their information. Its modular design allows for flexibility across various courses, while its AI-driven FAQ and keyword analysis help identify common student challenges.
Despite these advantages, potential risks must be considered, including over-reliance on AI and the accuracy of generated responses. While the LLM’s capabilities and limitations with regard to answering students’ questions are not the focus of this work, advanced LLM prompting strategies, such as Chain-of-Thought [17], and the use of a larger LLM, would be worthwhile improvements.
Furthermore, as students are aware that INSIGHT monitors their questions, they may ask questions to another LLM outside of INSIGHT. This can create a biased view for the teaching staff and skew their perception of students’ knowledge. To encourage the adoption of INSIGHT, students have to notice a positive impact on their learning experience using this tool, and thus, teaching staff need to actively use the new insights to improve student-teacher interactions. The teaching staff also has to communicate to the students when to use INSIGHT, as it is complementary to the staff and not a replacement. Certain questions are best answered by the teaching staff through direct, in-person interaction with students. Also note that INSIGHT fits within Hasselt University’s approach of interactive classes with teaching assistants, but may need to be revised to accommodate other class settings.
The next step in our research is to validate the usefulness of INSIGHT. To achieve this, we are applying for ethical approval to introduce INSIGHT as a teaching tool in, for example, the math summer courses and the consecutive mathematics courses at Hasselt University. Future research opportunities lie in integrating knowledge tracing to support advanced personalization while providing more granular data on the uptake of specific skills to the teaching staff.
4 Conclusion
INSIGHT is our attempt to exploit the potential of AI-driven educational tools to enhance both student learning and teaching staff support. INSIGHT’s goal is to help teaching staff identify knowledge gaps, personalize face-to-face support, and maintain student engagement by providing a modular, data-driven approach to analyzing student interactions. Its emphasis on privacy and transparency ensures that students can benefit from AI assistance without blindly giving away their data. In the next steps of our research, we will test INSIGHT’s effectiveness in real-world courses to validate our approach.
{credits}
4.0.1 Acknowledgements
This work was supported by the Special Research Fund (BOF) of Hasselt University (BOF24OWB28 and BOF23OWB31).
[…]
The authors have no competing interests to declare that are relevant to the content of this work.